# When Geoscience Meets Foundation Models: Towards General Geoscience Artificial Intelligence System

### v3

@@ II-C Potential Applications
Figure 3: Illustration of potential applications of Geoscience foundation model (GFM). On the left side, Earth science is depicted as encompassing four prominent disciplines: solid Earth science, atmospheric science, planetary science, and marine science. On the right side, various observation methods currently utilized in Earth science are represented, including space-based, air-based, ground-based, and ocean-based data collection. These observation methods provide multi-modal and multi-dimensional data that contribute to the field of Earth science. In the middle lies the GFM, which has the capability to integrate, correlate, and analyze the diverse geoscience data obtained from different observation means. By extracting higher-level features and representations from the data, GFMs enable enhanced understanding and insights. Once trained, this model can dynamically perform user-specified tasks in real-time, providing valuable support for Earth science research and applications.
@@
The geoscience field encompasses a diverse range of data types, each fulfilling a distinct role in research and analysis. These data types are pivotal in unraveling the mysteries of our planet and informing various scientific inquiries. As illustrated in Figure 3, the diverse array of observational modalities (space-air-ground-ocean) employed in geoscience research each provide unique affordances for characterizing the Earth system. Space-based data collection provides a global perspective, offering comprehensive coverage of the Earth’s surface including vegetation and ecosystems, the atmosphere and oceans. Air-based data collection serves as a bridge between space and terrestrial observations, enabling localized, high-resolution data acquisition. Ground-based data collection involves direct, in-situ measurements, enhancing our comprehension of geographical or geological occurrences. Additionally, ocean-based data collection plays a crucial role in evaluating marine and climatic conditions.
@@ II-C4 Ocean
III Recent Advances in Geoscience Foundation Models
With the applications of large foundation models for general purposes, the explorations of foundation models for remote sensing-based Geoscience tasks have drawn a lot of attention from the research community. In this subsection, we will review the recent advances in GFMs, including the key techniques of constructing GFMs, the summary of existing foundation models from the view of large language models (LLMs), large vision models (LVMs), and large language-vision models (LLVMs). Table I summarizes the existing foundation models, including data type, pre-training, adaptation, applications, and key features.
III-A Key Techniques of Constructing Geoscience Foundation Models
@@
As shown in Figure 4, pre-training and adaptation are two critical stages in the development and application of foundation models, where the former involves training a model on a large, diverse dataset to learn general-purpose representations, and the latter involves fine-tuning the pre-trained model on a specific task or dataset. The pre-training stage is particularly important for foundation models, as it enables the model to learn general-purpose representations that can be adapted to a wide range of tasks. The adaptation stage is equally important, as it allows the model to be fine-tuned for specific tasks or datasets, thereby improving its performance on those tasks.
@@
Figure 4: Key Techniques of Constructing Geoscience Foundation Models, Pre-training (a-d) and Fine-tuning (e-g).. a. Supervised Pre-training, training with large labeled data. b. Generative Pre-training, training by reconstructing data. c. Contrastive Pre-training, training by contrastive learning using similarity. d. Hybrid Pre-training, training with labeled and unlabeled data. e. Model Fine-tuning, fine-tuning GFMs using domain specific datasets. f. Prompt Engineering, fine-tuning GFMs using prompts. g. Prompt Tuning, fine-tuning GFMs using learned prompts.
III-A1 Pre-training
@@
Supervised pre-training
Supervised pre-training involves training a model on a large, labeled dataset. The model learns to predict the labels of the data, and the weights of the model are updated to minimize the prediction error. Supervised pre-training is particularly useful when there is a large, labeled dataset available for pre-training.
Self-supervised pre-training
Self-supervised pre-training involves training a model on a large, unlabeled dataset. The model learns to predict the missing parts of the data, and the weights of the model are updated to minimize the prediction error. Self-supervised pre-training is particularly useful when there is a large, unlabeled dataset available for pre-training. Generative pre-training trains a model to generate data that is similar to the input data by generating or reconstructing input data, while contrastive pre-training trains a model to distinguish between similar and dissimilar data.
Hybrid pre-training
Hybrid pre-training involves combining supervised and self-supervised pre-training. The model is first pre-trained on a large, labeled dataset using supervised pre-training, and then pre-trained on a large, unlabeled dataset using self-supervised pre-training. Hybrid pre-training is particularly useful when there is a large, labeled dataset available for pre-training, and a large, unlabeled dataset available for pre-training.
III-A2 Adaptation

### v4

@@ II-C Application Domain
Figure 3: Illustration of application domain of Geoscience foundation model (GFM). On the left side, Earth science is depicted as encompassing four prominent disciplines: solid Earth science, atmospheric science, planetary science, and marine science. On the right side, various observation methods currently utilized in Earth science are represented, including space-based, air-based, ground-based, and ocean-based data collection. These observation methods provide multi-modal and multi-dimensional data that contribute to the field of Earth science. In the middle lies the GFM, which has the capability to integrate, correlate, and analyze the diverse geoscience data obtained from different observation means. By extracting higher-level features and representations from the data, GFMs enable enhanced understanding and insights. Once trained, this model can dynamically perform user-specified tasks in real-time, providing valuable support for Earth science research and applications.
@@
The geoscience field encompasses a diverse range of data types, each fulfilling a distinct role in research and analysis. These data types are pivotal in unraveling the mysteries of our planet and informing various scientific inquiries. As illustrated in Fig. 3, the diverse array of observational modalities (space-air-ground-ocean) employed in geoscience research each provide unique affordances for characterizing the Earth system. Space-based data collection provides a global perspective, offering comprehensive coverage of the Earth’s surface including vegetation and ecosystems, the atmosphere and oceans. Air-based data collection serves as a bridge between space and terrestrial observations, enabling localized, high-resolution data acquisition. Ground-based data collection involves direct, in-situ measurements, enhancing our comprehension of geographical or geological occurrences. Additionally, ocean-based data collection plays a crucial role in evaluating marine and climatic conditions.
@@ II-C4 Ocean
III Key Techniques of Constructing Geoscience Foundation Models
@@
In this section, we discuss the key techniques for constructing geoscience foundation models (GFMs), including transformers, pre-training and adaptation. As shown in Fig. 4, pre-training and adaptation are two critical stages in the development and application of foundation models, where the former involves training a model on a large, diverse dataset to learn general-purpose representations, and the latter involves fine-tuning the pre-trained model on a specific task or dataset. The pre-training stage is particularly important for foundation models, as it enables the model to learn general-purpose representations that can be adapted to a wide range of tasks. The adaptation stage is equally important, as it allows the model to be fine-tuned for specific tasks or datasets, thereby improving its performance on those tasks.
@@
Figure 4: Key Techniques of Constructing Geoscience Foundation Models, including Pre-training (a-d) and Adaptation (e-i). a. Supervised Pre-training, training with large labeled data. b. Generative Pre-training, training by reconstructing data. c. Contrastive Pre-training, training by contrastive learning using similarity. d. Hybrid Pre-training, training with labeled and unlabeled data. e. Fine-tuning, fine-tuning GFMs using domain specific datasets. f. Prompt Tuning, fine-tuning GFMs using prompts. g. Adapter Tuning, fine-tuning GFMs using additional trainable parameters. h. Parameter Tuning, fine-tuning GFMs by modifying specific parameters. i. Remapping Tuning, fine-tuning GFMs by knowledge transfer.
III-A Transformers
The Transformer architecture, introduced by Vaswani et al. [78], represents a significant leap forward in neural network design for natural language processing (NLP). At its core, the Transformer’s power stems from its self-attention mechanisms, which enable effective modeling of long-range dependencies and contextual relationships in language [79]. This approach addresses limitations observed in traditional sequential models such as recurrent neural networks (RNNs) and convolutional neural networks (CNNs), which often struggle with processing long sequences [80, 81]. By calculating correlations between each token and all others in the input sequence, the Transformer mitigates challenges faced by conventional models, making it a versatile tool for various NLP applications. For more details, please refer to the related literatures [82, 83, 84].
III-A1 Vanilla Transformer
The Transformer’s architecture is primarily composed of an encoder and a decoder, each featuring multi-head self-attention mechanisms and position-wise feed-forward networks [78]. In the encoder, input sequences are projected into multiple vector spaces, undergoing a series of linear transformations, attention computations, and concatenation operations to generate context-aware output vectors. The decoder uses these context vectors to generate the output sequence, with an additional attention layer to focus on relevant parts of the input, similar to the attention mechanism in Seq2Seq models [85]. This structure, as illustrated in Fig. 5, allows the Transformer to process extended text sequences with equal consideration given to all positions, offering advantages over recurrent architectures in terms of parallelization and long-range dependency modeling [86, 87, 88, 89].
Figure 5: The Transformer - model architecture [78].
III-A2 Visual Transformer
The Transformer architecture has revolutionized NLP with remarkable success. Its impact has since expanded into computer vision, demonstrating impressive capabilities across various tasks such as image recognition, classification, and segmentation. The Transformer has emerged as a versatile and scalable framework, offering significant advantages in training efficiency compared to traditional methods. Its flexibility allows for implementation as a standalone architecture or in combination with CNNs to achieve superior results in visual tasks. Typical examples include the vision transformer (ViT) [90], which has garnered widespread attention for its convolution-free design and exceptional ability to model long-range dependencies.
ViT architecture [90] is illustrated in Fig. 6. Initially, the input image is segmented into $16\times 16$ patches, which are then flattened and linearly projected. These processed patches, combined with positional encodings, are fed into the Transformer encoder. The encoded features are subsequently passed through an MLP head for classification. The transformer encoder primarily utilizes the positional encoding structure proposed in the original Transformer model. ViT represents a pioneering application of the Transformer architecture to computer vision tasks [89]. Its convolution-free design and exceptional ability to model long-range dependencies have garnered significant attention from researchers. Building upon ViT’s foundation, several variants have emerged, including DeiT [91], TNT [92], and PVT [93]. These iterations have further enhanced performance by incorporating advanced techniques such as knowledge distillation, hierarchical token mixing, and pyramid-style vision transformers.
Figure 6: The framework of ViT [90].
III-A3 Vision-Language Transformer
Vision-language transformers bridge the gap between computer vision and natural language processing. These models leverage the powerful Transformer architecture to simultaneously process and understand both visual and textual information, enabling a wide range of multimodal tasks. By learning joint representations of images and text, vision-language transformers can perform complex operations such as visual question answering, image captioning, and cross-modal retrieval with remarkable efficiency and accuracy. Most representative models include VisualBERT [94], Uniter [95], and OSCAR [96]. These models integrate BERT for text encoding and Faster R-CNN for object detection, enabling the joint processing of visual and textual inputs through multiple transformer layers.
VisualBERT [94] pioneered the implicit alignment of textual elements with image regions using self-attention mechanisms. This model integrated BERT [97] for NLP with a pretrained Faster-RCNN [98] for object detection. The approach treated the original text and image features extracted from object proposals as unordered input tokens, processing them jointly through multiple Transformer layers to capture complex cross-modal associations (as shown in Fig. 7). Following VisualBERT’s innovative approach, several VLMs emerged, including Uniter [95], OSCAR [96], and InterBert [99]. These models similarly employed BERT as their text encoder and Faster-RCNN for object proposal generation, further advancing the modeling of vision-language interactions.
Figure 7: The architecture of VisualBERT[94].
III-B Pre-training
@@
III-B1 Supervised pre-training
Supervised learning has been a cornerstone in the development of deep learning models, driving success in both vision and language tasks. In computer vision, models including AlexNet [101], VGG [102], and ResNet [103] achieved breakthrough performance through supervised pre-training on large-scale datasets such as ImageNet [104], JFT-300M [105], and Kinetics [106]. Similarly, in natural language processing, early transformer-based models such as BERT [97], RoBERTa [107], and ALBERT [108] utilized supervised pre-training on tasks such as masked language modeling and sentence ordering prediction. As the field progressed, researchers explored larger-scale and more diverse pre-training approaches, such as Gato’s [109] multi-modal, multi-task learning, and multi-label learning in vision models scaling up to 22 billion parameters [110]. In the biomedical domain, BioBERT [111] specializes in medical text analysis, while ELECTRA [112] employs a novel discriminator-based approach for efficient language model pre-training. VideoBERT [113] extends to multimodal learning for video understanding. This evolution in pre-training strategies highlights the ongoing quest for more efficient and generalizable learning paradigms in AI, balancing the need for task-specific knowledge with broader applicability to avoid overfitting.
III-B2 Self-supervised pre-training
Self-supervised learning (SSL) has become a cornerstone in the development of foundation models (FMs), enabling them to acquire rich features and representations from vast amounts of unlabeled data [114], [115], [116]. This approach, often referred to as the ”dark matter” of intelligence, circumvents the limitations of supervised learning by leveraging a ”semiautomatic” process to derive ”labels” directly from the data, significantly reducing time and resource costs [115]. SSL has achieved remarkable success across various domains of deep learning, particularly in natural language processing with models like BERT [97] and GPT-3 [117], and in computer vision with MAE [118] and DINOv2 [119]. The SSL workflow typically involves defining a pretext task based on unlabeled inputs to generate meaningful representations [114], [116], [120], [121]. These pretext tasks, which can be prediction-based, context-based, or generation-based, form the basis of the pretraining phase where the objective function is learned from unlabeled data [120]. The resulting representations are then transferred as initial weights for fine-tuning on specific downstream tasks. SSL methods can be broadly categorized into three approaches based on their pretext tasks: generative, contrastive, and predictive [114], [122]. This pretraining-finetuning paradigm has proven highly effective, allowing FMs to leverage the power of unlabeled data before being adapted to specific tasks with labeled data, often matching or surpassing purely supervised approaches in performance.
{forest}
[…]
training, fill=gray!60, text=white, rounded corners [Supervised, fill=blue!80, text=white, rounded corners [CNNs [AlexNet [101]] [VGG [102]] [ResNet [103] ] ] [Transformers [BERT [97]] [RoBERTa [107]] ] ] [SSL, fill=blue!80, text=white, rounded corners [Generative [GAN [BigGAN [123]] [SRGAN [124]] ] [Autoencoder [VAE [125]] [MAE [118]] ] ] [Contrastive [Negative sampling [SimCLR [126]] [MoCo [127]] ] [Clustering [DeepCluster [128]] [SwAV [129]] ] [Knowledge
distillation [BYOL [130]] [DINO [131]] ] [Redundancy
reduction [Barlow Twins [132]] ] ] [Predictive [Spatial [CFN [133]] [RotNet [134]] ] [Spectral [Colorful Image
Colorization [135]] ] ] ] [Hybrid, fill=blue!80, text=white, rounded corners [ELECTRA [112]] [DALL-E [136]] ] ]
Figure 8: Taxonomy of the pre-training methods with examples.
Generative pre-training
Generative methods in self-supervised learning focus on learning representations through the process of reconstructing or generating input data. These approaches seek to model the underlying data distribution, thereby capturing intrinsic statistical properties and dependencies within the data [137]. By doing so, generative methods can implicitly extract meaningful features and structures without the need for explicit labels. This capability makes them particularly valuable in scenarios where labeled data is scarce or expensive to obtain. Common implementations of generative methods often employ sophisticated models such as autoencoders, which compress and then reconstruct input data, or generative adversarial networks (GANs), which pit a generator against a discriminator in a competitive learning process.
Autoencoders form the foundation of many generative methods, consisting of an encoder network that maps input data to a latent space representation and a decoder network that reconstructs the data from this latent space. This structure has spawned several variants, notably the variational autoencoder (VAE) [125], which incorporates probabilistic modeling by assuming a prior distribution $P(Z)$ over the latent space and modeling the conditional distribution $P(X|Z)$. In VAEs, the encoder approximates the posterior distribution $P(Z|X)$ by inferring the model $Q(Z|X)$. Recent applications of generative methods include information recovery tasks such as inpainting [133], where a network’s context encoder restores missing pixel values based on surrounding context. This concept has evolved into masked autoencoders (MAEs) [118], which apply random masks to input image patches and reconstruct the missing pixels. The process of accurately reconstructing each pixel requires the model to develop a comprehensive understanding of the various objects and components in the image, resulting in learned feature representations that prove valuable for a wide range of downstream tasks.
GANs [137] represent a powerful approach in generative methods, consisting of two competing networks: a generator $G:Z\rightarrow X$ that synthesizes fake samples from random noise in the latent space, and a discriminator $D:X\rightarrow[0,1]$ that attempts to distinguish between real and fake samples. This adversarial setup has proven effective for tasks such as image generation and super-resolution. Notable advancements include BigGAN [123], which introduces innovations like a large-scale GAN architecture, conditional batch normalization, category conditioning, and orthogonal regularization, enabling the generation of high-quality, diverse, and category-specific images. In the domain of image enhancement, SRGAN [124] has demonstrated remarkable capabilities in super-resolution tasks, learning from a large dataset of downsampled images to recover high-resolution textures and effectively upscale low-resolution inputs. These applications showcase the potential of GANs not just for generating new images, but also for enhancing existing ones, pushing the boundaries of synthetic image generation and opening up numerous possibilities in image processing and computer vision tasks.
Contrastive pre-training
Contrastive learning methods aim to learn representations by maximizing similarity between related samples and minimizing it between unrelated ones. This approach uses a contrastive objective function to bring similar samples closer in feature space while separating dissimilar ones. The key advantage of contrastive learning is its ability to extract meaningful representations without explicit labels, leveraging the data’s inherent structure. These methods typically fall into four categories: negative sampling, clustering, knowledge distillation, and redundancy reduction.
Negative sampling in contrastive learning involves creating dissimilar examples to anchor samples, aiming to differentiate positive and negative instances in the representation space. This approach helps learn valuable features while avoiding model collapse. Notable implementations include SimCLR [126], which uses a CNN feature extractor with diverse data augmentations, and the MoCo series [127]. MoCo employs a momentum-based update strategy and compares samples against a large queue of negative examples. MoCo and its variants (MoCov2 [138], MoCov3 [139]) have shown strong performance in self-supervised learning tasks, demonstrating the effectiveness of negative sampling in developing robust and discriminative representations. These methods have significantly advanced the field of contrastive learning, providing powerful tools for learning from unlabeled data in various computer vision tasks.
The clustering approach in contrastive learning groups similar samples unsupervised. DeepCluster [128] exemplifies this method by first clustering images and then training a CNN to recognize these assignments. Another notable example is SwAV [129], which introduces the concept of grouping similar representations. SwAV generates multiple views of an image through data augmentation, clusters these views based on similarity, and encourages similar representations within clusters while differentiating between clusters. These methods demonstrate how unsupervised clustering can effectively create meaningful groupings in the feature space, enabling the model to learn rich representations without explicit labels. By leveraging the inherent structure in the data, clustering approaches in contrastive learning contribute to developing more robust and adaptable models for various computer vision tasks.
Knowledge distillation in self-supervised learning (SSL) transfers knowledge from a high-performing teacher model to a student model. Notable examples include BYOL [130] and the DINO series [131], [119]. BYOL predicts representations of augmented views, using an online network as the student and a momentum-updated target network as the teacher. DINO combines transformer architecture with consistency maximization between image views, where one view predicts the teacher network’s output for another. DINOv2 enhances this approach by adding a clustering objective, encouraging semantically similar samples to cluster in the learned representation. These methods demonstrate how knowledge distillation can effectively leverage pretrained models to guide the learning of robust, fine-grained representations in SSL, advancing the field of unsupervised visual representation learning.
Redundancy reduction in contrastive learning aims to minimize redundant information in learned representations, focusing on capturing essential and discriminative features. A notable example is the Barlow twins model [132], which optimizes the cross-correlation matrix of representations. This approach maximizes cross-correlation between different views of the input while minimizing diagonal elements, effectively encouraging the model to encode useful, non-redundant information. By balancing the capture of statistical dependencies with redundancy reduction, this method enables the learning of more efficient and informative representations. This approach contributes to the development of more compact and effective feature representations, potentially improving model performance and generalization across various computer vision tasks.
Predictive pre-training
Predictive methods in self-supervised learning aim to learn useful representations through various prediction tasks, encompassing both spatial and spectral aspects of images. Spatial-based approaches include relative position prediction, where CNNs are trained to predict the positions of randomly sampled image patches [140]; image puzzles, which involve reconstructing images from shuffled patches to learn object part relationships and shapes; geometric transformation recognition, such as RotNet [134], which identifies rotation angles of images; and counting tasks that train models to enumerate visual primitives. Spectral-based methods, exemplified by image colorization (IC) tasks [135], focus on automatically adding realistic colors to grayscale images. These diverse approaches encourage models to learn rich, contextual representations by solving pretext tasks that leverage the inherent structure and properties of images. For instance, the Jigsaw puzzle approach in CFN [140] trains networks to understand spatial arrangements of object parts, while IC tasks help models learn complex color distributions and object relationships. By engaging with these challenging tasks, models develop robust feature representations that prove valuable for downstream applications such as classification, detection, and segmentation, demonstrating the effectiveness of predictive methods in unsupervised visual learning.
III-B3 Hybrid pre-training
Hybrid pre-training approaches have significantly advanced the development of large AI models across multiple domains, combining diverse learning techniques to leverage their strengths and address limitations. In large language models (LLMs), examples include ELECTRA’s [112] generator-discriminator setup and BART’s [141] combination of bidirectional encoding and autoregressive decoding. Large vision models (LVMs) often pair contrastive learning with supervised tasks, as seen in SimCLR [126] and MAE [118]. Large vision-language models (LVLMs) such as CLIP [142] and DALL-E [136] utilize cross-modal contrastive learning, combining vision and language inputs. These approaches typically involve self-supervised pre-training on large unlabeled datasets followed by supervised fine-tuning. By integrating multiple pre-training objectives and architectures, hybrid approaches enable models to capture diverse aspects of their domains – from language understanding and generation to visual feature extraction and cross-modal reasoning – resulting in improved performance across a wide range of downstream tasks.
III-C Adaptation